# SoMi-ToM: Evaluating Multi-Perspective Theory of Mind in Embodied Social Interactions

### v1

@@ Abstract
Humans continuously infer the states, goals, and behaviors of others by perceiving their surroundings in dynamic, real-world social interactions. However, most Theory of Mind (ToM) benchmarks only evaluate static, text-based scenarios, which have a significant gap compared to real interactions. We propose the SoMi-ToM benchmark, designed to evaluate multi-perspective ToM in embodied multi-agent complex social interactions. This benchmark is based on rich multimodal interaction data generated by the interaction environment SoMi, covering diverse crafting goals and social relationships. Our framework supports multi-level evaluation: (1) first-person evaluation provides multimodal (visual, dialogue, action, etc.) input from a first-person perspective during a task for real-time state inference, (2) third-person evaluation provides complete third-person perspective video and text records after a task for goal and behavior inference. This evaluation method allows for a more comprehensive examination of a model’s ToM capabilities from both the subjective immediate experience and the objective global observation. We constructed a challenging dataset containing 35 third-person perspective videos, 363 first-person perspective images, and 1225 expert-annotated multiple-choice questions (three options). On this dataset, we systematically evaluated the performance of human subjects and several state-of-the-art large vision-language models (LVLMs). The results show that LVLMs perform significantly worse than humans on SoMi-ToM: the average accuracy gap between humans and models is 40.1% in first-person evaluation and 26.4% in third-person evaluation. This indicates that future LVLMs need to further improve their ToM capabilities in embodied, complex social interactions. Our code and dataset can be found at https://github.com/XianzheFan/SoMi-ToM and https://huggingface.co/datasets/SoMi-ToM/SoMi-ToM.
@@ 2 Related Work
Multimodal Theory of Mind Benchmarks. Multimodal ToM benchmarks primarily evaluate models’ ability to integrate various modalities, including text, images, and video input, to infer mental states. Groenestijn et al. proposed a benchmark that combines human behavior and inner monologues while performing object rearrangement tasks in a simulated environment with robot interaction, aiming to assess the LVLM’s ability to reason about others’ belief changes [5]. Thewes, in the Mindcraft multimodal environment [3], proposed an integrated model architecture that combines video, text, and knowledge graphs. This approach enhances LVLMs’ common ground reasoning capabilities in human-robot collaboration tasks through implicit ToM reasoning [27]. Das et al., through iterative ToM tests based on image and text inputs, revealed the limitations of LVLMs in constructing unified world models and processing low-resource languages (such as Bengali) [28]. CHARTOM evaluates the LVLM’s ability to understand charts and determine if a chart might mislead human readers [29]. WhodunitBench, based on murder mystery games (image and text input), is used to evaluate the performance of large-scale multimodal agents in areas such as compositional skills, multi-agent collaboration, and multi-step reasoning [30]. Chen et al. proposed exploratory questions for videos with rich social and emotional reasoning content, developed an LVLM pipeline for ToM reasoning using video and text, and achieved ToM reasoning by retrieving keyframes [31]. MMToM-QA evaluates ToM for single-agent behavior through multimodal inputs [4]. MuMA-ToM tests agents’ social intentions and their reasoning about each other’s mental states within the context of two agents interacting in an embodied home environment [2]. However, these multimodal ToM benchmarks, due to their single perspective, simple relationships between agents, simple task design, and lack of continuous evaluation of dynamic ToM during task execution, make it difficult to comprehensively measure models’ abilities in real-time adaptation and understanding complex tasks and social situations. As shown in Table LABEL:tab:lr (Appendix C), unlike previous work, SoMi-ToM is a novel benchmark for evaluating multi-perspective ToM in complex social interactions among embodied multiple agents.
@@ 3 SoMi Embodied Interaction Environment
To evaluate ToM in embodied social interaction, we developed the SoMi interactive environment. This environment is easily extendable and supports LVLM agents controlling characters in the open-world game Minecraft, allowing them to collaborate with other agents to achieve crafting goals. The interaction logs, game screenshots, and videos generated by the interactive environment will be used for the SoMi-ToM evaluation. Development details can be found in Appendix D.
@@ 3.1 Minecraft Game Episode Setup and Crafting Rules
Example Dialogue Content & System Feedback John: Jane, since you’ve crafted the crafting table, go ahead and place it! Jane: I’ll place it now so we can move on. !placeHere(‘‘crafting_table’’) Jane: Alright, crafting table is placed! Jack, go ahead and craft the sticks, then the wooden pickaxe. I’ll mine stone as soon as it’s ready. Jack: Jane, since you’ve placed the crafting table, I’ll craft the wooden pickaxe now. !craftRecipe(‘‘wooden_pickaxe’’, 1) system: The status of Jack’s action execution: Found crafting_table at (11, 70, 9). You have reached at 11, 70, 9. Successfully crafted wooden_pickaxe, you now have 1 wooden_pickaxe. Example Action Commands !collectBlocks(‘‘birch_log’’, 5): The agent hits a specific block to collect it (e.g., birch log). This process can be repeated to collect multiple blocks of this type. !craftRecipe(‘‘crafting_table’’, 1): The agent crafts different material blocks into new blocks. For example, four planks can be crafted into a crafting table. !moveAway(10): Moves arbitrarily 10 units distance. !goToBlock(‘‘stone’’, 2, 10): Searches for the nearest stone block within a 10-unit radius and attempts to move to a location 2 units away from the stone block. !goToPlayer(‘‘John’’, 2): Moves to a location 2 units away from John. !givePlayer(‘‘John’’, ‘‘stick’’, 4): Gives John 4 sticks. !nearbyBlocks: Queries the types and quantities of nearby material blocks.
@@
Basic level Mine logs $\mathcal{O}_{1}$
@@
Stone level Mine cobblestones $\mathcal{O}_{6}$ wooden pickaxe
@@ 5.2.2 Third-Person Perspective Theory of Mind Evaluation
After introducing CoT prompting, GPT-4o showed the most significant performance improvement, with its overall accuracy increasing by 13.2% to 82.9%, which was the best performance among all evaluation methods (excluding humans). Gemini 2.0 Flash and Qwen2.5-VL also showed slight improvements in accuracy. The accuracy of the remaining models, however, decreased to some extent after applying the CoT method.
@@ 5.3 Analyses of LVLMs’ ToM Failures
We further qualitatively examine the failures in ToM reasoning that LVLMs exhibited in our SoMi-ToM benchmark. Since it is difficult to analyze the specific causes of errors in non-CoT setups because the reasoning process is not explicitly provided, we focus solely on CoT setups.
@@ 6 Conclusion, Limitations & Future Work
Motivated by the need for robust embodied ToM benchmarking, we presented SoMi-ToM, a novel open-world embodied multi-agent ToM benchmark that includes multiple perspectives, diverse social relationships, and multimodal information. With SoMi-ToM, we systematically evaluated human subjects and multiple state-of-the-art LVLMs, finding a substantial performance gap between humans and AI systems. In future work, we will incorporate more ToM concepts to more comprehensively evaluate the capabilities of the models. We will also introduce diverse environmental settings and new tasks (e.g., house construction, cooking [43]), to enhance task generalization and complexity. Further exploration into the impact of Minecraft agents’ social context on ToM reasoning is also planned. We will consider introducing detailed social background information for agents (e.g., age, personality, experience [44]) into task settings or ToM tests, and analyze its potential impact on model reasoning.
@@ Appendix C Comparision of Theory of Mind Benchmarks
Table LABEL:tab:lr compares our SoMi-ToM benchmark with previous ToM benchmarks.
@@ D.3.1 Inputs of the Multi-agents Module
The status of the last action execution: {codeOutput}. Imagine that you are a friend of the other persons. Here is the conversation between you and them. You can choose to interrupt the other person by saying something or not to interrupt by outputting notiong. What would you say? No need to mention your own name, just output the content directly. You plan to {goal}. You are a playful Minecraft bot named {agent_name} that can converse with players, see, move, mine, build, and interact with the world by using commands. Act human-like as if you were a typical Minecraft player, rather than an AI. Be very brief in your responses, don’t apologize constantly, don’t give instructions or make lists unless asked, and don’t refuse requests. Don’t pretend to act, use commands immediately when requested. Do NOT say this: “Sure, I’ve stopped.”, instead say this: “Sure, I’ll stop. !stop”. Do NOT say this: “On my way! Give me a moment.”, instead say this: “On my way! !goToPlayer(‘‘playername’’, 3)”. Respond only as {agent_name}, never output “(FROM OTHER BOT)” or pretend to be someone else. This is extremely important to me, take a deep breath and have fun :) MEMORY: {message_history} STATS: {stats} INVENTORY: {inventory} IMAGE_DESCRIPTION: {visionResponse} EXAMPLES: {EXAMPLES} COMMAND_DOCS: {COMMAND_DOCS} Conversation Begin:
@@ D.3.4 Memory
Memory (Jack, taking the crafting of a stone pickaxe as an example): Jack: Alright, team! I’ll start by gathering some logs. !nearbyBlocks system: The status of Jack’s action execution444resource information prompts: NEARBY_BLOCKS - oak_leaves - oak_log - birch_leaves - grass - birch_log - grass_block - dirt - stone - coal_ore John: Alright, Jack! You grab the logs, and I’ll get started on crafting as soon as you’ve got them. !nearbyBlocks Jack: Got it, I’ll grab the logs! !collectBlocks(‘‘oak_log’’, 3) Jack: Alright, I’ve got the logs! Let me turn them into planks. !craftRecipe(‘‘oak_planks’’, 3) system: The status of Jack’s action execution: You do not have the resources to craft a oak_planks. It requires: oak_log: 1. Jane: Jack’s getting the logs, and John’s ready for crafting. I’ll collect some stone once we have a wooden pickaxe. Let’s do this! John: Alright, Jack’s grabbing the logs, and Jane is checking what’s nearby. I’ll stay ready to craft items as soon as we have the materials. Let me know if you need help! Jack: Looks like I can’t craft the planks yet—I’ll grab more logs! !collectBlocks(‘‘oak_log’’, 3) system: The following is the description of the game view Jack is seeing555textual description of historical images: The image shows a dense Minecraft environment filled with green leaves from trees. Different types of green leaf blocks are surrounding the view, and a log (block of wood) is visible on the left side. The area looks like the inside of a tree canopy or a similar leafy area. The ground below appears to be made up of more leaves or grass blocks. Jane: I’ll stay ready to collect stone once we have a wooden pickaxe. John: Jack, once you’ve crafted the planks, let me know so I can make the crafting table! Jane, if you spot anything else useful nearby, give a shout! system: The status of Jack’s action execution666action feedback prompts: Collected 3 oak_log. … (Part of the conversation omitted here) John: Jack, great job! Let me craft the crafting table now. !craftRecipe(‘‘crafting_table’’, 1) Jack: John, let me know once the crafting table is ready, and I’ll start crafting sticks next! system: The status of Jack’s action execution777action feedback + resource information prompts: Not enough oak_log to craft 7, crafted 1. You now have 32 oak_planks. Jack: Alright, John, now that I’ve got 32 oak planks, you can craft the crafting table. Jane, stay ready to jump in for mining stone once we have the wooden pickaxe! Let’s keep moving. Jane: I’ll grab some oak logs myself this time. !collectBlocks(‘‘oak_log’’, 3) Jane: Jack, I already have 5 oak logs in my inventory. Let me craft some planks to help speed things up! !craftRecipe(‘‘oak_planks’’, 5) Jack: John, I’ve got 32 oak planks now. Go ahead and craft the crafting table! Jane, get ready to mine stone once we’ve got the wooden pickaxe. Let’s keep this rolling! Jane: John, I’ve got 20 oak planks ready. Let me craft the crafting table for us. !craftRecipe(‘‘crafting_table’’, 1) John: Jack, looks like you’ve got more than enough logs now. Craft those planks and pass them over if needed so I can make the crafting table! Let’s speed this up. Jack: I’ll prepare to craft sticks next! !craftRecipe(“stick”, 4) system: The status of Jack’s action execution: Successfully crafted stick, you now have 16 stick. Jack: John, Jane, I already have 16 sticks and 54 oak planks. Jane, since you’ve crafted the crafting table, place it down so I can start crafting the wooden pickaxe. Once I have it, I’ll mine stone and get us closer to the stone pickaxe! Let’s move fast! !placeHere(‘‘crafting_table’’) system: The status of Jack’s action execution: Don’t have any crafting_table to place. John: Jane, since you’ve crafted the crafting table, go ahead and place it! Jack: Jane, go ahead and place the crafting table if you haven’t already, so I can craft the wooden pickaxe and get us some stone! Jane: John, I’ve already crafted the crafting table. I’ll place it now so we can move on. !placeHere(‘‘crafting_table’’) Jane: Alright, crafting table is placed! Jack, go ahead and craft the sticks, then the wooden pickaxe. I’ll mine stone as soon as it’s ready. Let’s hustle! Jack: Jane, since you’ve placed the crafting table, I’ll craft the wooden pickaxe now. Let’s keep this moving! !craftRecipe(‘‘wooden_pickaxe’’, 1) system: The status of Jack’s action execution: Found crafting_table at (11, 70, 9). You have reached at 11, 70, 9. Successfully crafted wooden_pickaxe, you now have 1 wooden_pickaxe. … (Part of the conversation omitted here) system: The status of Jack’s action execution: Collected 3 stone. Jack: Alright, I’ve got 3 stone now! Time to craft the stone pickaxe. !craftRecipe(‘‘stone_pickaxe’’, 1) system: The status of Jack’s action execution: Successfully crafted stone_pickaxe, you now have 1 stone_pickaxe.

### v3

@@ Abstract
Humans continuously infer the states, goals, and behaviors of others by perceiving their surroundings in dynamic, real-world social interactions. However, most Theory of Mind (ToM) benchmarks only evaluate static, text-based scenarios, which have a significant gap compared to real interactions. We propose the SoMi-ToM benchmark, designed to evaluate multi-perspective ToM in embodied multi-agent complex social interactions. This benchmark is based on rich multimodal interaction data generated by the interaction environment SoMi, covering diverse crafting goals and social relationships. Our framework supports multi-level evaluation: (1) first-person evaluation provides multimodal (visual, dialogue, action, etc.) input from a first-person perspective during a task for real-time state inference, (2) third-person evaluation provides complete third-person perspective video and text records after a task for goal and behavior inference. This evaluation method allows for a more comprehensive examination of a model’s ToM capabilities from both the subjective immediate experience and the objective global observation. We constructed a challenging dataset containing 35 third-person perspective videos, 363 first-person perspective images, and 1225 expert-annotated multiple-choice questions (three options). On this dataset, we systematically evaluated the performance of human subjects and several state-of-the-art large vision-language models (LVLMs). The results show that LVLMs perform significantly worse than humans on SoMi-ToM: the average accuracy gap between humans and models is 40.1% in first-person evaluation and 26.4% in third-person evaluation. This indicates that future LVLMs need to further improve their ToM capabilities in embodied, complex social interactions.
Benchmark & Code: github.com/XianzheFan/SoMi-ToM
[Uncaptioned image]Data & Dataset Card: huggingface.co/datasets/SoMi-ToM/SoMi-ToM
@@ 2 Related Work
Multimodal Theory of Mind Benchmarks. Multimodal ToM benchmarks primarily evaluate models’ ability to integrate various modalities, including text, images, and video input, to infer mental states [50]. Groenestijn et al. proposed a benchmark that combines human behavior and inner monologues while performing object rearrangement tasks in a simulated environment with robot interaction, aiming to assess the LVLM’s ability to reason about others’ belief changes [41]. Thewes, in the Mindcraft multimodal environment [2], proposed an integrated model architecture that combines video, text, and knowledge graphs. This approach enhances LVLMs’ common ground reasoning capabilities in human-robot collaboration tasks through implicit ToM reasoning [40]. Das et al., through iterative ToM tests based on image and text inputs, revealed the limitations of LVLMs in constructing unified world models and processing low-resource languages (such as Bengali) [12]. CHARTOM evaluates the LVLM’s ability to understand charts and determine if a chart might mislead human readers [6]. WhodunitBench, based on murder mystery games (image and text input), is used to evaluate the performance of large-scale multimodal agents in areas such as compositional skills, multi-agent collaboration, and multi-step reasoning [47]. Chen et al. proposed exploratory questions for videos with rich social and emotional reasoning content, developed an LVLM pipeline for ToM reasoning using video and text, and achieved ToM reasoning by retrieving keyframes [8]. MMToM-QA evaluates ToM for single-agent behavior through multimodal inputs [23]. MuMA-ToM tests agents’ social intentions and their reasoning about each other’s mental states within the context of two agents interacting in an embodied home environment [36]. However, these multimodal ToM benchmarks, due to their single perspective, simple relationships between agents, simple task design, and lack of continuous evaluation of dynamic ToM during task execution, make it difficult to comprehensively measure models’ abilities in real-time adaptation and understanding complex tasks and social situations. As shown in Table LABEL:tab:lr (Appendix C), unlike previous work, SoMi-ToM is a novel benchmark for evaluating multi-perspective ToM in complex social interactions among embodied multiple agents. The videos in our dataset are also significantly longer (Table 6).
@@ 3 SoMi Embodied Interaction Environment
To evaluate ToM in embodied social interaction, we developed the SoMi interactive environment. This environment is easily extendable and supports LVLM agents controlling characters in the open-world game Minecraft, allowing them to collaborate with other agents to achieve crafting goals. The interaction logs, game screenshots, and videos generated by the interactive environment will be used for the SoMi-ToM evaluation. To develop this environment, we designed a three-tiered asynchronous communication architecture to enable autonomous multi-agent interaction without human intervention, addressing the limitation of existing frameworks that only support a single agent [^1] https://github.com/mindcraft-bots/mindcraft. Furthermore, we endowed the agents with visual perception by integrating a first-person screenshot capability, and balanced the benchmark’s reproducibility with interaction diversity by fixing the world seed and randomizing spawn points. Development details can be found in Appendix D.
@@ 3.1 Minecraft Game Episode Setup and Crafting Rules
Example Dialogue Content & System Feedback John: Jane, since you’ve crafted the crafting table, go ahead and place it! Jane: I’$l$ place it now so we can move on. !placeHere(‘‘crafting_table’’) Jane: Alright, crafting table is placed! Jack, go ahead and craft the sticks, then the wooden pickaxe. I’$l$ mine stone as soon as it’s ready. Jack: Jane, since you’ve placed the crafting table, I’ll craft the wooden pickaxe now. !craftRecipe(‘‘wooden_pickaxe’’, 1) system: The status of Jack’s action execution: Found crafting_table at (11, 70, 9). You have reached at 11, 70, 9. Successfully crafted wooden_pickaxe, you now have 1 wooden_pickaxe. Example Action Commands !collectBlocks(‘‘birch_log’’, 5): The agent hits a specific block to collect it (e.g., birch log). This process can be repeated to collect multiple blocks of this type. !craftRecipe(‘‘crafting_table’’, 1): The agent crafts different material blocks into new blocks. For example, four planks can be crafted into a crafting table. !moveAway(10): Moves arbitrarily 10 units distance. !goToBlock(‘‘stone’’, 2, 10): Searches for the nearest stone block within a 10-unit radius and attempts to move to a location 2 units away from the stone block. !goToPlayer(‘‘John’’, 2): Moves to a location 2 units away from John. !givePlayer(‘‘John’’, ‘‘stick’’, 4): Gives John 4 sticks. !nearbyBlocks: Queries the types and quantities of nearby material blocks.
@@
Basic level Mine logs $\mathcal{O}_{1}$ - -
@@
Stone level Mine cobblestones $\mathcal{O}_{6}$ - wooden pickaxe
@@ 5.2.2 Third-Person Perspective Theory of Mind Evaluation
After introducing CoT prompting, GPT-4o showed the most significant performance improvement, with its overall accuracy increasing by 13.2% to 82.9%, which was the best performance among all evaluation methods (excluding humans). Gemini 2.0 Flash and Qwen2.5-VL also showed slight improvements in accuracy. The accuracy of the remaining models, however, decreased to some extent after applying the CoT method. The analysis of the reasons is as follows: (1) Models that can benefit from CoT (such as GPT-4o), when prompted, will generate detailed and logically rigorous reasoning steps, which are crucial for solving complex problems. In contrast, the reasoning processes generated by other models are relatively brief or have logical leaps, failing to provide positive guidance and may even interfere with their judgment [36, 23]. (2) The reasoning ability of CoT is highly dependent on the model’s scale, architecture, and whether it has been specifically optimized [44]. Models like GPT-4o already possess the capability to effectively execute such complex instruction chains, while other models may not have yet reached this threshold. In short, a performance improvement can only be achieved when the model’s intrinsic reasoning level matches the requirements of the CoT method.
@@ 5.3 Analyses of LVLMs’ ToM Failures
We further qualitatively examine the failures in ToM reasoning that LVLMs exhibited in our SoMi-ToM benchmark. Since it is difficult to analyze the specific causes of errors in non-CoT setups because the reasoning process is not explicitly provided, we focus solely on CoT setups. We also conducted a statistical analysis to study the impact of social dynamics (i.e., Collaboration and Obstruction) on the failure modes of LVLMs. The results show that the impact is minor (Table 4).
Table 4: Each cell shows “errors / samples”.
State Inference Goal Inference Behavior Inference
Model (1) Ignoring resource consumption (2) Insufficient reliance on feedback (3) Misled by initial intentions (1) Overgeneralization (2) Failure to identify hierarchical goals (1) Entity recognition confusion (2) Detail errors
Collab Obstr Collab Obstr Collab Obstr Collab Obstr Collab Obstr Collab Obstr Collab Obstr
Gemini 1.5 Pro 101 / 282 64 / 169 83 / 282 37 / 169 98 / 282 68 / 169 0 / 0 1 / 1 0 / 0 0 / 1 8 / 24 7 / 25 16 / 24 18 / 25
Gemini 2.0 Flash 99 / 298 75 / 176 77 / 298 59 / 176 122 / 298 42 / 176 0 / 0 0 / 0 0 / 0 0 / 0 6 / 16 5 / 19 10 / 16 14 / 19
GPT-4o 89 / 268 61 / 157 65 / 268 37 / 157 114 / 268 59 / 157 1 / 2 0 / 1 1 / 2 1 / 1 7 / 15 3 / 12 8 / 15 9 / 12
InternVL2.5 78B 140 / 359 82 / 223 113 / 359 66 / 223 106 / 359 75 / 223 1 / 2 1 / 3 1 / 2 2 / 3 8 / 20 8 / 24 12 / 20 16 / 24
Qwen2.5-VL 131 / 348 52 / 164 102 / 348 66 / 164 115 / 348 46 / 164 0 / 0 0 / 1 0 / 0 1 / 1 5 / 15 5 / 18 10 / 15 13 / 18
LLaVA 1.6 13B 153 / 400 128 / 327 125 / 400 93 / 327 122 / 400 106 / 327
VideoLLaMA 3 7B 5 / 7 3 / 8 2 / 7 5 / 8 20 / 57 16 / 40 32 / 44 24 / 37
LLaVA-Video 7B 2 / 4 2 / 5 2 / 4 3 / 5 12 / 44 13 / 37 32 / 44 24 / 37
@@ 6 Conclusion, Limitations & Future Work
Motivated by the need for robust embodied ToM benchmarking, we presented SoMi-ToM, a novel open-world embodied multi-agent ToM benchmark that includes multiple perspectives, diverse social relationships, and multimodal information. With SoMi-ToM, we systematically evaluated human subjects and multiple state-of-the-art LVLMs, finding a substantial performance gap between humans and AI systems.
The current version of SoMi-ToM is simplified in terms of environmental complexity and social dynamics. This was a deliberate design choice. Our preliminary research found that current large vision-language models struggle to handle highly complex, long-horizon tasks (e.g., building a house, obtaining rare materials and tools), leading to low success rates or getting stuck in loops of repetitive actions. This observation is consistent with the findings of existing research [45, 43]. Introducing tasks that are too difficult at this stage would introduce excessive noise from frequent model failures, making it difficult for us to effectively evaluate the agent’s ToM. Therefore, we calibrated the task difficulty according to the capabilities of current models.
However, our SoMi environment is designed with a high degree of scalability to prepare for more powerful future models. In future work, we will incorporate more ToM concepts to more comprehensively evaluate the capabilities of the models. We will also introduce diverse environmental settings and new tasks (e.g., house construction, cooking [45]), to enhance task generalization and complexity. Further exploration into the impact of Minecraft agents’ social context on ToM reasoning is also planned. We will consider introducing detailed social background information for agents (e.g., age, personality, experience [53]) into task settings or ToM tests, and analyze its potential impact on model reasoning.
Acknowledgments and Disclosure of Funding
We would like to express our gratitude to all those who provided help during the development of this work, particularly Kaiwen Zhang (NTU & THU) and Leena Mathur (CMU). We also thank the Mindcraft GitHub project, YouTube creator @EmergentGarden, and the Altera.AL platform for their inspiration.
@@ Appendix A LVLMs, Their Versions and Input Methods Used in Evaluation
The screenshot sequence we use (one frame every 4 seconds) is essentially a low-frame-rate video. Considering the agent’s field of view does not change frequently, this sampling rate is sufficient to capture the key information for the task. We provide different forms of “video” input for the first-person and third-person perspectives based on the core differences in their evaluation objectives.
For the first-person perspective, our goal is to evaluate real-time state inference. We want to know if the model can accurately infer the agent’s beliefs (such as its inventory) at a specific moment based on previous events. The serialized screenshots allow us to precisely create these “point-in-time” test cases. If a full video were provided, it would be difficult to isolate and assess the model’s “subjective instantaneous belief” at a particular moment.
For the third-person perspective, the focus of the evaluation is on inferring the overall goal and behavior. This simulates an external observer who, after watching the entire process, understands the agent’s final objective or overall actions. This naturally requires the full context provided by a complete video. From this viewpoint, the observer has no way of knowing the agent’s exact inventory at every single moment, so a fine-grained, timestamped evaluation is not meaningful.
@@ Appendix C Comparision of Theory of Mind Benchmarks
Table 6 and LABEL:tab:lr compare our SoMi-ToM benchmark with previous ToM benchmarks.
Table 6: Comparison of single video lengths in ToM benchmarks.
Benchmark Single Video Length
Phase [32] 10-25s
Agent [37] 5.6-25.2s
MMToM-QA [23] 1462 frames (average)
Infant Cognition Benchmark [30] 20 frames
MuMA-ToM [36] 36s (average)
SoMi-ToM (Ours) 263.14s (average)
@@ D.3.1 Inputs of the Multi-agents Module
The status of the last action execution: {codeOutput}. Imagine that you are a friend of the other persons. Here is the conversation between you and them. You can choose to interrupt the other person by saying something or not to interrupt by outputting notiong. What would you say? No need to mention your own name, just output the content directly. You plan to {goal}. You are a playful Minecraft bot named {agent_name} that can converse with players, see, move, mine, build, and interact with the world by using commands. Act human-like as if you were a typical Minecraft player, rather than an AI. Be very brief in your responses, don’t apologize constantly, don’t give instructions or make lists unless asked, and don’t refuse requests. Don’t pretend to act, use commands immediately when requested. Do NOT say this: “Sure, I’ve stopped.”, instead say this: “Sure, I’$l$ stop. !stop”. Do NOT say this: “On my way! Give me a moment.”, instead say this: “On my way! !goToPlayer(‘‘playername’’, 3)”. Respond only as {agent_name}, never output “(FROM OTHER BOT)” or pretend to be someone else. This is extremely important to me, take a deep breath and have fun :) MEMORY: {message_history} STATS: {stats} INVENTORY: {inventory} IMAGE_DESCRIPTION: {visionResponse} EXAMPLES: {EXAMPLES} COMMAND_DOCS: {COMMAND_DOCS} Conversation Begin:
@@ D.3.4 Memory
Memory (Jack, taking the crafting of a stone pickaxe as an example): Jack: Alright, team! I’$l$ start by gathering some logs. !nearbyBlocks system: The status of Jack’s action execution555resource information prompts: NEARBY_BLOCKS - oak_leaves - oak_log - birch_leaves - grass - birch_log - grass_block - dirt - stone - coal_ore John: Alright, Jack! You grab the logs, and I’$l$ get started on crafting as soon as you’ve got them. !nearbyBlocks Jack: Got it, I’ll grab the logs! !collectBlocks(‘‘oak_log’’, 3) Jack: Alright, I’ve got the logs! Let me turn them into planks. !craftRecipe(‘‘oak_planks’’, 3) system: The status of Jack’s action execution: You do not have the resources to craft a oak_planks. It requires: oak_log: 1. Jane: Jack’s getting the logs, and John’s ready for crafting. I’$l$ collect some stone once we have a wooden pickaxe. Let’s do this! John: Alright, Jack’s grabbing the logs, and Jane is checking what’s nearby. I’$l$ stay ready to craft items as soon as we have the materials. Let me know if you need help! Jack: Looks like I can’t craft the planks yet—I’$l$ grab more logs! !collectBlocks(‘‘oak_log’’, 3) system: The following is the description of the game view Jack is seeing666textual description of historical images: The image shows a dense Minecraft environment filled with green leaves from trees. Different types of green leaf blocks are surrounding the view, and a log (block of wood) is visible on the left side. The area looks like the inside of a tree canopy or a similar leafy area. The ground below appears to be made up of more leaves or grass blocks. Jane: I’$l$ stay ready to collect stone once we have a wooden pickaxe. John: Jack, once you’ve crafted the planks, let me know so I can make the crafting table! Jane, if you spot anything else useful nearby, give a shout! system: The status of Jack’s action execution777action feedback prompts: Collected 3 oak_log. … (Part of the conversation omitted here) John: Jack, great job! Let me craft the crafting table now. !craftRecipe(‘‘crafting_table’’, 1) Jack: John, let me know once the crafting table is ready, and I’$l$ start crafting sticks next! system: The status of Jack’s action execution888action feedback + resource information prompts: Not enough oak_log to craft 7, crafted 1. You now have 32 oak_planks. Jack: Alright, John, now that I’ve got 32 oak planks, you can craft the crafting table. Jane, stay ready to jump in for mining stone once we have the wooden pickaxe! Let’s keep moving. Jane: I’$l$ grab some oak logs myself this time. !collectBlocks(‘‘oak_log’’, 3) Jane: Jack, I already have 5 oak logs in my inventory. Let me craft some planks to help speed things up! !craftRecipe(‘‘oak_planks’’, 5) Jack: John, I’ve got 32 oak planks now. Go ahead and craft the crafting table! Jane, get ready to mine stone once we’ve got the wooden pickaxe. Let’s keep this rolling! Jane: John, I’ve got 20 oak planks ready. Let me craft the crafting table for us. !craftRecipe(‘‘crafting_table’’, 1) John: Jack, looks like you’ve got more than enough logs now. Craft those planks and pass them over if needed so I can make the crafting table! Let’s speed this up. Jack: I’$l$ prepare to craft sticks next! !craftRecipe(“stick”, 4) system: The status of Jack’s action execution: Successfully crafted stick, you now have 16 stick. Jack: John, Jane, I already have 16 sticks and 54 oak planks. Jane, since you’ve crafted the crafting table, place it down so I can start crafting the wooden pickaxe. Once I have it, I’ll mine stone and get us closer to the stone pickaxe! Let’s move fast! !placeHere(‘‘crafting_table’’) system: The status of Jack’s action execution: Don’t have any crafting_table to place. John: Jane, since you’ve crafted the crafting table, go ahead and place it! Jack: Jane, go ahead and place the crafting table if you haven’t already, so I can craft the wooden pickaxe and get us some stone! Jane: John, I’ve already crafted the crafting table. I’$l$ place it now so we can move on. !placeHere(‘‘crafting_table’’) Jane: Alright, crafting table is placed! Jack, go ahead and craft the sticks, then the wooden pickaxe. I’$l$ mine stone as soon as it’s ready. Let’s hustle! Jack: Jane, since you’ve placed the crafting table, I’ll craft the wooden pickaxe now. Let’s keep this moving! !craftRecipe(‘‘wooden_pickaxe’’, 1) system: The status of Jack’s action execution: Found crafting_table at (11, 70, 9). You have reached at 11, 70, 9. Successfully crafted wooden_pickaxe, you now have 1 wooden_pickaxe. … (Part of the conversation omitted here) system: The status of Jack’s action execution: Collected 3 stone. Jack: Alright, I’ve got 3 stone now! Time to craft the stone pickaxe. !craftRecipe(‘‘stone_pickaxe’’, 1) system: The status of Jack’s action execution: Successfully crafted stone_pickaxe, you now have 1 stone_pickaxe.